# Is This Collection Worth My LLM’s Time? Automatically Measuring Information Potential in Text Corpora

### v2

@@ 4.4 Qualitative Analysis of High-Value Information
  3. 3.
Additional examples of these patterns can be found in Table 3.
5 Conclusion
This work introduces a novel and efficient approach for evaluating the information potential of text collections for large language models. Our key contributions include:
  1. 1.
    Efficient Evaluation Pipeline: We present a systematic approach combining automated MCQ generation, sophisticated filtering mechanisms, and comparative evaluation to assess information potential without requiring model training or fine-tuning.
  2. 2.
    Empirical Validation: Our results validate the method’s effectiveness by demonstrating alignment with intuitive expectations across both dataset types and model scales. The information potential increases from synthetic baseline (0.125) to Wikipedia articles (0.136-0.142) to EPFL manuscripts (0.229) to Venetian historical records (0.267), while larger models (GPT-4o) consistently show lower information potential than smaller ones (Llama 70B), suggesting better knowledge retention.
  3. 3.
    Qualitative Framework: We propose a taxonomy of high-value information types (technical terminology, novel research findings, and complex relationships), providing deeper insights into the nature of valuable information sources for LLMs.
The successful application of our method to both PhD manuscripts and the Venetian historical collection highlights its versatility across specialized knowledge domains, reinforcing its potential as a practical tool for evaluating diverse information sources—from contemporary academic research to rare historical records. This cross-domain applicability underscores the importance of identifying areas where large language models (LLMs) struggle to retrieve information, as these “gray zones” may correspond to gaps in pretraining data. As isolating the dataset’s intrinsic contribution would require costly fine-tuning experiments, directly correlating information gain with missing pretraining data remains challenging. Nonetheless, recognizing these limitations is crucial. Just as students who are aware of their weakest subjects can better prepare for an exam—perhaps even bringing a well-structured cheat sheet if allowed—models that can identify and incorporate these gray zones as input in domain-specific tasks may achieve greater proficiency in handling specialized knowledge.
@@ 6 Limitations
Practical Considerations. Our approach, while effective, presents several methodological limitations that suggest directions for future research. First, the current pipeline’s generation-first approach, where questions are created before applying filtering criteria, could be enhanced by incorporating context-answer alignment and distractor plausibility directly into the generation process. Such an integrated approach could improve question quality while reducing computational overhead.
Second, the scope of our evaluation faces an inherent challenge: identifying datasets that are definitively outside the training distribution of large language models. While PhD manuscripts offer a reasonable proxy for novel content, the growing scale and opacity of training corpora make it increasingly difficult to verify whether any given dataset was truly unseen. This limitation underscores the urgency of developing methods to assess and prioritize valuable textual resources before they are absorbed into future training data.
Finally, our method focuses exclusively on selection rather than integration. While our results highlight where information gaps likely exist, they do not resolve how selected content might be optimally incorporated into LLMs. Emerging work, such as SKILL (Moiseev et al., 2022), proposes promising directions for efficient knowledge injection, while others (Liu et al., 2024) caution against potential interference with previously learned representations. Addressing these tensions remains an important challenge for future research.
On Human Evaluation and Pre-training Experiments. This study also omits human evaluation and pre-training-based validation. This is a deliberate design choice: our goal is to develop a lightweight, scalable method that can be used when such resource-intensive steps are impractical. Rather than replacing these approaches, our pipeline is designed as a screening mechanism. Importantly, our results exhibit strong alignment with expected knowledge gaps across datasets (e.g., low information potential for synthetic text, moderate for Wikipedia, high for novel PhD and historical records), providing empirical support for the method’s validity. Furthermore, we integrate multiple safeguards—including MCQ quality filters, position-debiased evaluation, and multi-model testing—to enhance robustness. While future work may explore downstream fine-tuning or human validation, we argue that our approach already provides reliable, cost-effective insight for strategic dataset selection.
Broader Considerations. While our approach requires digital text for MCQ generation, it offers a significantly more efficient alternative to full model training or fine-tuning. This enables strategic sampling approaches where representative portions of larger collections can be evaluated to inform broader digitization decisions. For instance, assessing a few chapters can inform decisions about entire book collections, making the method particularly valuable for resource-constrained scenarios. This sampling-based approach could revolutionize how institutions prioritize their digitization efforts, allowing for data-driven decisions about resource allocation in preservation projects.
More broadly, our effort to categorize high value information sheds light on fundamental challenges in distinguishing between information that is unknown to an LLM versus information it fails to retrieve. This raises important questions about the relationship between knowledge possession and practical competence in AI systems. When an LLM consistently fails to demonstrate knowledge in a specific domain, the distinction between these cases may become less relevant from a practical perspective. This observation has broader implications for how we conceptualize and evaluate knowledge in AI systems, suggesting that performance-based metrics can also bring complementary analysis to theoretical attempts that map internal knowledge representations.
@@ A.2.1 Positional Bias
ABCD001010101020202020303030304040404050505050PercentageBaselineWikipedia-EFLEPFL
@@
ABCD001010101020202020303030304040404050505050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With Context
@@
ABCD001010101020202020303030304040404050505050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With ContextBaseline - No ContextBaseline - With Context
Figure 6: Distribution of Correct Answer Letter Prediction for EPFL, Wikipedia, and Baseline MCQ Datasets Evaluated on Llama 70B After Positional Bias Mitigation.
@@ A.2.2 Information Potential
Figure 7 shows the models’ performance with an emphasis on the positional bias mitigation strategy. As every MCQ is evaluated four times, this allows the analysis of a model’s consistency. It is clearly visible that the models become less consistent without context in datasets with higher information potential while this trend is less pronounced with context. Figure 8 shows the performance of Llama 70B along with the information potential across the datasets and cutoff percentiles.
(a) Llama 70B EPFL
@@
(d) Llama 70B Wikipedia-EPFL
@@
001010101020202020303030304040404050505050606060600.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Cutoff Percentile ValuesFraction of Correct ResponsesBaselineWikipedia-EPFLEPFLWC 4xNC 4xIP
@@
A.2.3 Qualitative Analysis of EPFL MCQ Dataset

### v3

@@ 5 Conclusion
    Evaluation pipeline: Automated MCQ generation, filtering, and position-debiased testing provide a scalable measure of collection value.
@@
    Empirical validation: Results align with expectations—IP rises from synthetic baseline (0.125) to Wikipedia (0.136–0.180) to PhD manuscripts (0.229) and Venetian records (0.265). Larger models (GPT-4o) show lower IP than smaller ones (Llama 3 70B), consistent with stronger prior knowledge.
By highlighting where LLMs benefit most from external context, our method offers a practical screening tool for prioritizing digitization and integration of specialized corpora, from recent academic research to rare historical sources.
@@ 6 Limitations
Practical Considerations. Our pipeline generates MCQs before filtering; integrating quality checks directly into generation could improve efficiency. A further challenge is verifying whether corpora truly lie outside LLM training data, given opaque and ever-growing pretraining sets. Finally, our method identifies promising collections but does not address how to integrate them into models without disrupting existing knowledge (Moiseev et al., 2022; Liu et al., 2024).
Evaluation Scope. This study omits human annotation and pretraining experiments to keep the method lightweight. Instead, we rely on safeguards such as similarity filtering, position-debiased testing, and cross-model evaluation. While fine-tuning or human validation could strengthen findings, our results already align with intuitive knowledge gaps (e.g., low IP for synthetic, high IP for specialized corpora).
Broader Perspective. The method requires digitized text but supports sampling-based evaluation—e.g., testing a few chapters before digitizing entire collections—making it useful for resource-constrained projects. More generally, IP highlights areas where models lack or fail to retrieve knowledge, raising broader questions about the line between knowledge possession and competence, and offering a practical complement to theoretical analyses of internal representations.
@@ A.2.1 Positional Bias
ABCD010102020303040405050PercentageBaselineWikipedia-EFLEPFL
@@
ABCD010102020303040405050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With Context
@@
ABCD010102020303040405050PercentageEPFL - No ContextEPFL - With ContextWikipedia-EPFL - No ContextWikipedia-EPFL - With ContextBaseline - No ContextBaseline - With Context
Figure 5: Distribution of Correct Answer Letter Prediction for EPFL, Wikipedia, and Baseline MCQ Datasets Evaluated on Llama 3 70B After Positional Bias Mitigation.
@@ A.2.2 Information Potential
Figure 6 shows the models’ performance with an emphasis on the positional bias mitigation strategy. As every MCQ is evaluated four times, this allows the analysis of a model’s consistency. It is clearly visible that the models become less consistent without context in datasets with higher information potential while this trend is less pronounced with context. Figure 7 shows the performance of Llama 3 70B along with the information potential across the datasets and cutoff percentiles.
(a) Llama 3 70B EPFL
@@
(d) Llama 3 70B Wikipedia-EPFL
@@
01010202030304040505060600.20.20.40.40.60.60.80.811Cutoff Percentile ValuesFraction of Correct ResponsesBaselineWikipedia-EPFLEPFLWC 4xNC 4xIP
@@
A.2.3 Qualitative Analysis of High-Value Information
MCQ (correct answer in italics) Relevant Context Passages
B) Both typically result from incorrect or incomplete mental models.
C) Both involve action reversals.
A) Repetition.
B) Omission.
C) Timing error.
D) Sequence error.
A) AFs rely on spectral analysis.
B) AFs use a frame-to-phoneme alignment.
C) AFs map phonemes to articulatory features.
D) AFs predict phonemes directly.
  1. 1.
  2. 2.
  3. 3.
Additional examples of these patterns can be found in Table 3.
A.2.4 Qualitative Analysis of EPFL MCQ Dataset